# Poincaré inequality and Quantitative De Giorgi method for hypoelliptic operators

### v2

@@ 1.5. Motivation
Kolmogorov equations appear in the theory of stochastic processes: equation (1.2) was studied by Kolmogorov in 1934 [11] when $A$ is the identity matrix to understand the “time-integrated Brownian motion”. The equation is then the Kolmogorov forward equation for the process
@@
Kolmogorov obtained explicit formula for its fundamental solution in [11] which inspired the seminal work [9]. Integrating twice in time the Brownian motion would lead to equation (1.4) with $\kappa=2$ and $A$ the identity matrix. Due to this connection to stochastic processes, several mathematical models involving linear and nonlinear Kolmogorov type equations have also appeared in finance, and in particular equations of type (1.4) appear in various models for pricing of path-dependent financial instruments. For example the equation for $P=P(S,A,t)$
@@ 1.6. Strategy of proof
The strategy to obtain a Poincaré inequality for the local case with one commutator in [8] can be summarised as follows. The first step is to mollify the characteristic function of $Q^{-}$, by adding an error term
$\int_{Q^{+}}\left(f-\langle f\rangle_{Q^{-}}\right)_{+}{\,\mathrm{d}}z\leq$ $\int_{z_{+}\in Q^{+}}\left(\int_{z_{-}\in Q^{-}}\left[f(z_{+})-f(z_{-})\right]\varphi_{\varepsilon}(w,y){\,\mathrm{d}}z_{-}\right)_{+}{\,\mathrm{d}}z_{+}$
@@
with $0\leq\varphi_{\varepsilon}\leq 1$ smooth, $\varepsilon>0$, $z_{+}=(x,v,t)\in Q^{+}$ and $z_{-}=(y,w,s)\in Q^{-}$. The second step consists in constructing piecewise affine trajectories following the two vector fields $\mathcal{T}:=\partial_{t}+v\cdot\nabla_{x}$ and $\nabla_{v}$ to connect any $z_{+}\in Q^{+}$ with any $z_{-}\in Q^{-}$:
@@
Note here that a positive time gap between the two cylinders $Q^{-}$ and $Q^{+}$ is used to make sure the intermediate velocity $(x-y)/(t-s)$ remains bounded. We then write $f(z_{+})-f(z_{-})$ as an integral along the chosen trajectory and use ${\mathcal{T}}f+{\mathcal{A}}^{*}{\mathcal{A}}f\leq 0$:
$\int_{Q^{-}}\left[f(z_{+})-f(z_{-})\right]\varphi_{\varepsilon}(z_{-}){\,\mathrm{d}}z_{-}$ $\lesssim\int_{Q^{-}}\int_{\text{trajectory}}{\mathcal{T}}f(\cdots)\varphi_{\varepsilon}(w,y){\,\mathrm{d}}s{\,\mathrm{d}}z_{-}$
$\quad+\int_{Q^{-}}\int_{\text{trajectory}}\nabla_{v}f(\cdots)\varphi_{\varepsilon}(w,y){\,\mathrm{d}}s{\,\mathrm{d}}z_{-}$
$\lesssim\int_{Q^{-}}\int_{\text{trajectory}}\left[\nabla_{v}\cdot\left(A\nabla_{v}f\right)\right](\cdots)\varphi_{\varepsilon}(w,y){\,\mathrm{d}}s{\,\mathrm{d}}z_{-}$
$\quad+\int_{Q^{-}}\int_{\text{trajectory}}\nabla_{v}f(\cdots)\varphi_{\varepsilon}(w,y){\,\mathrm{d}}s{\,\mathrm{d}}z_{-}.$
@@
$(v,x,t)$ $\underset{\nabla_{x}}{\longrightarrow}(v,x+\varepsilon w,t)\underset{\nabla_{v}}{\longrightarrow}\left(\frac{x+\varepsilon w-y}{t-s},x+\varepsilon w,t\right)$
$\underset{{\mathcal{T}}}{\longrightarrow}\left(\frac{x+\varepsilon w-y}{t-s},y,s\right)\underset{\nabla_{v}}{\longrightarrow}(w,y,s).$
This sub-trajectory “noises” the future position variable with the past velocity. This produces a second error term of the form $\varepsilon^{\sigma}\|f\|_{L^{1}_{t,v}W^{\sigma,1}_{x}}$, which can, however, be controlled for $\varepsilon$ small by the integral regularity $L^{1}_{t,v}W^{\sigma,1}_{x}$ (which can be established for sub-solutions to the Kolmogorov equation, for a small $\sigma>0$). Finally it yields
@@
for a constant $C_{\varepsilon}>0$ depending on $\varepsilon>0$. The two additional errors weakens the Poincaré inequality, but the latter is shown to be sufficient for implementing the next steps of the De Giorgi theory in [8].
@@
(1.22) $\fint_{Q^{+}}\left[f(z_{+})-\langle f\rangle_{Q^{-}}\right]_{+}{\,\mathrm{d}}z_{+}$ $\leq\fint_{Q^{+}}\left(\fint_{Q^{0}}\left[f(z_{+})-f(z_{0})\right]\varphi_{\varepsilon}(v_{0},x_{0}){\,\mathrm{d}}z_{0}\right)_{+}{\,\mathrm{d}}z_{+}$
$\quad+\fint_{Q^{-}}\left(\fint_{Q^{0}}\left[f(z_{0})-f(z_{-})\right]\varphi_{\varepsilon}(v_{0},x_{0}){\,\mathrm{d}}z_{0}\right)_{+}{\,\mathrm{d}}z_{-}$
@@ 1.7. Outline
In Section 2 we provide the explicit construction of the trajectories connecting points in the future to points in the past. These trajectories are then used in Section 3 to prove the Poincaré inequality; they allow to estimate the $L^{1}$ norm in the future of the difference between $f$ and the past average of $f$, or in other words, the left hand side of (1.18). In Section 4, we prove the Harnack inequalities on the basis of this Poincaré inequality and deduce the Hölder regularity.
@@ 2.1. The base case
(2.1) $\begin{cases}{\frac{\mathrm{d}}{\mathrm{d}s}}X^{(0)}_{\pm}(s)=\mathsf{m}_{\pm}^{(0)}g_{0}^{\prime\prime}(s)+\mathsf{m}^{(1)}_{\pm}g_{1}^{\prime\prime}(s),\\[5.69054pt] {\frac{\mathrm{d}}{\mathrm{d}s}}X^{(1)}_{\pm}(s)=\delta_{\pm}X^{(0)}_{\pm}(s),\\[5.69054pt] {\frac{\mathrm{d}}{\mathrm{d}s}}T_{\pm}(s)=\delta_{\pm},\end{cases}$
@@
-$\nabla_{(0)}$$\nabla_{(0)}$${\mathcal{T}}$$\dot{\Gamma}^{+}(s)$$x^{(0)}$$t$$x^{(1)}$$\Gamma_{+}(0)$$\Gamma_{+}(1)$
Figure 3. Construction of the trajectories. The curve $\Gamma_{+}$ connects any point $\Gamma_{+}(0)=z_{+}\in Q^{+}$ to some intermediate point $\Gamma_{+}(1)=z_{0}\in Q^{0}$ along the vector fields ${\mathcal{T}}$p and $\nabla_{(0)}$.
@@
(2.2) $\begin{cases}X^{(0)}_{\pm}(s)=x^{(0)}_{\pm}+\mathsf{m}_{\pm}^{(0)}g_{0}^{\prime}(s)+\mathsf{m}^{(1)}_{\pm}g_{1}^{\prime}(s),\\[5.69054pt] X^{(1)}_{\pm}(s)=x^{(1)}_{\pm}+s\delta_{\pm}x^{(0)}_{\pm}+\delta_{\pm}\left[\mathsf{m}_{\pm}^{(0)}g_{0}(s)+\mathsf{m}^{(1)}_{\pm}g_{1}(s)\right],\\[5.69054pt] T_{\pm}(s)=st_{0}+(1-s)t_{\pm}.\end{cases}$
Let us denote $\text{{\bf 1}}_{d}$ the $d$-vector of $1$’s, $\text{{\bf 0}}_{d}$ the $d$-vector of 00’s, $\text{Id}_{d}$ the $d\times d$ identity matrix, $\mathbb{O}$ a zero block of arbitrary size, and
@@
${\mathsf{W}}_{\pm}^{\delta}(s):=\begin{pmatrix}\text{Id}_{d}&\mathbb{O}\\ \mathbb{O}&\delta_{\pm}\text{Id}_{d}\end{pmatrix}{\mathsf{W}}(s),\quad\mathsf{Y}_{\pm}=\begin{pmatrix}\mathsf{y}^{(0)}_{\pm}\\ \mathsf{y}^{(1)}_{\pm}\end{pmatrix}:=\begin{pmatrix}x^{(0)}_{0}-x^{(0)}_{\pm}\\ x^{(1)}_{0}-x^{(1)}_{\pm}-\delta_{\pm}x^{(0)}_{\pm}\end{pmatrix}.$
@@
$\begin{pmatrix}X^{(0)}_{\pm}(s)\\ X^{(1)}_{\pm}(s)\end{pmatrix}={\mathsf{W}}_{\pm}^{\delta}(s)\mathsf{M}_{\pm}+\begin{pmatrix}\text{Id}_{d}&\mathbb{O}\\ \delta_{\pm}s\text{Id}_{d}&\text{Id}_{d}\end{pmatrix}\begin{pmatrix}x_{\pm}^{(0)}\\ x_{\pm}^{(1)}\end{pmatrix}.$
@@
$\begin{pmatrix}X^{(0)}_{\pm}(s)\\ X^{(1)}_{\pm}(s)\end{pmatrix}$ $={\mathsf{W}}_{\pm}^{\delta}(s){\mathsf{W}}_{\pm}^{\delta}(1)^{-1}\mathsf{Y}_{\pm}+\begin{pmatrix}\text{Id}_{d}&\mathbb{O}\\ \delta_{\pm}s\text{Id}_{d}&\text{Id}_{d}\end{pmatrix}\begin{pmatrix}x_{\pm}^{(0)}\\ x_{\pm}^{(1)}\end{pmatrix}$
@@
$\quad-W^{\delta}_{\pm}(s)\big{[}W^{\delta}_{\pm}(1)\big{]}^{-1}\begin{pmatrix}x^{(0)}_{\pm}\\ x^{(1)}_{\pm}+\delta_{\pm}x^{(0)}_{\pm}\end{pmatrix}+\begin{pmatrix}\text{Id}_{d}&\mathbb{O}\\ \delta_{\pm}s\text{Id}_{d}&\text{Id}_{d}\end{pmatrix}\begin{pmatrix}x_{\pm}^{(0)}\\ x_{\pm}^{(1)}\end{pmatrix}$
@@
${\mathsf{W}}_{\pm}^{\delta}(1):=\begin{pmatrix}\frac{\text{Id}_{d}}{(1+\alpha_{0})}&\frac{\text{Id}_{d}}{(1+\alpha_{1})}\\ \frac{\delta_{\pm}\text{Id}_{d}}{(1+\alpha_{0})(2+\alpha_{0})}&\frac{\delta_{\pm}\text{Id}_{d}}{(1+\alpha_{1})(2+\alpha_{1})}\end{pmatrix},$
${\mathsf{W}}_{\pm}^{\delta}(s)^{-1}:=\frac{1}{\delta_{\pm}\det{\mathsf{W}}(s)}\begin{pmatrix}\frac{\delta_{\pm}s^{2+\alpha_{1}}\text{Id}_{d}}{(1+\alpha_{1})(2+\alpha_{1})}&\frac{-s^{1+\alpha_{1}}\text{Id}_{d}}{(1+\alpha_{1})}\\ \frac{-\delta_{\pm}s^{(2+\alpha_{0})}\text{Id}_{d}}{(1+\alpha_{0})(2+\alpha_{0})}&\frac{s^{1+\alpha_{0}}\text{Id}_{d}}{(1+\alpha_{0})}\end{pmatrix},$
@@ 2.2. The general case
(2.5) $\begin{cases}{\frac{\mathrm{d}}{\mathrm{d}s}}X^{(0)}_{\pm}(s)=\sum_{i=0}^{\kappa}\mathsf{m}_{\pm}^{(i)}g_{i}^{(\kappa+1)}(s),\\[5.69054pt] {\frac{\mathrm{d}}{\mathrm{d}s}}X^{(1)}_{\pm}(s)=\delta_{\pm}{\mathsf{B}}_{1}X^{(0)}_{\pm}(s),\\[5.69054pt] \quad\vdots\\[5.69054pt] {\frac{\mathrm{d}}{\mathrm{d}s}}X^{(\kappa)}_{\pm}(s)=\delta_{\pm}{\mathsf{B}}_{\kappa}X^{(\kappa-1)}_{\pm}(s),\\[5.69054pt] {\frac{\mathrm{d}}{\mathrm{d}s}}T_{\pm}(s)=\delta_{\pm}\end{cases}$
@@
(2.6) $\begin{cases}X^{(0)}_{\pm}(s)=x^{(0)}_{\pm}+\sum_{i=0}^{\kappa}\mathsf{m}_{\pm}^{(i)}g_{i}^{(\kappa)}(s),\\[5.69054pt] X^{(1)}_{\pm}(s)=x^{(1)}_{\pm}+s\delta_{\pm}\tilde{\mathsf{B}}_{1,1}x^{(0)}_{\pm}+\delta_{\pm}\left[\sum_{i=0}^{\kappa}\tilde{\mathsf{B}}_{1,1}\mathsf{m}_{\pm}^{(i)}g_{i}^{(\kappa-1)}(s)\right],\\[5.69054pt] X^{(2)}_{\pm}(s)=x^{(2)}_{\pm}+s\delta_{\pm}\tilde{\mathsf{B}}_{2,2}x^{(1)}_{\pm}+\frac{(s\delta_{\pm})^{2}}{2}\tilde{\mathsf{B}}_{2,1}x^{(0)}_{\pm}+\delta_{\pm}^{2}\left[\sum_{i=0}^{\kappa}\tilde{\mathsf{B}}_{2,1}\mathsf{m}_{\pm}^{(i)}g_{i}^{(\kappa-2)}(s)\right],\\[5.69054pt] \quad\vdots\\[5.69054pt] X^{(\kappa)}_{\pm}(s)=x_{\pm}^{(\kappa)}+\sum_{i=1}^{\kappa}\frac{(s\delta_{\pm})^{i}}{i!}\tilde{\mathsf{B}}_{\kappa,\kappa-i+1}x^{(\kappa-i)}_{\pm}+\delta_{\pm}^{\kappa}\left[\sum_{i=0}^{\kappa}\tilde{\mathsf{B}}_{\kappa,1}\mathsf{m}_{\pm}^{(i)}g_{i}(s)\right],\\[5.69054pt] T_{\pm}(s)=st_{0}+(1-s)t_{\pm}\end{cases}$
where $\tilde{\mathsf{B}}_{i,j}:={\mathsf{B}}_{i}{\mathsf{B}}_{i-1}\cdots{\mathsf{B}}_{j}$ for $1\leq j\leq i$, which is a $d_{i}\times d_{j-1}$ block matrix. We denote
@@
(2.12) $\nabla_{(0)}\left(\Phi_{\pm}^{s}\right)^{-1}={\mathsf{W}}_{\pm}^{\delta}(1)\big{[}{\mathsf{W}}_{\pm}^{\delta}(s)\big{]}^{-1}\begin{pmatrix}\hbox{Id}_{d_{0}}\\ \mathbb{O}_{d_{1}}\\ \vdots\\ \mathbb{O}_{d_{\kappa}}\end{pmatrix}.$
@@
Consider the $(\kappa+1)\times(\kappa+1)$ matrix $P(s):=(s^{1+\alpha_{j}+i}/p_{i,j})_{i,j=0}^{\kappa}$ with $p_{i,j}:=(1+\alpha_{j})\cdots(1+i+\alpha_{j})$ the matrix [^1] The notation $P\otimes\text{Id}_{d_{0}}$ is the classical tensor product between the matrices $P$ and $\text{Id}_{d_{0}}$. $W:=P\otimes\text{Id}_{d_{0}}$. Then their determinants are given by

### v3

@@ 1.5. Motivation
Kolmogorov equations appear in the theory of stochastic processes: equation (1.3) was studied by Kolmogorov in 1934 [12], when $A$ is the identity matrix, to understand the “time-integrated Brownian motion”. The equation is then the Kolmogorov forward equation for the process
@@
Kolmogorov obtained the explicit formula for its fundamental solution in [12] which inspired the seminal work [10]. Integrating twice in time the Brownian motion would lead to equation (1.5) with $\kappa=2$ and $A$ the identity matrix. Due to this connection to stochastic processes, several mathematical models involving linear and nonlinear Kolmogorov type equations have also appeared in finance, and in particular equations of type (1.5) appear in various models for pricing of path-dependent financial instruments. For example the equation for $P=P(S,A,t)$
@@ 1.6. Strategy of proof
The strategy to obtain a Poincaré inequality for the local case when $\kappa=1$ in [8] can be summarised as follows. The first step is to mollify the characteristic function of $Q^{-}$, by adding an error term
$\int_{Q^{+}}\left(f-\langle f\rangle_{Q^{-}}\right)_{+}{\,\mathrm{d}}z\leq$ $\int_{z_{+}\in Q^{+}}\left(\int_{z_{-}\in Q^{-}}\left[f(z_{+})-f(z_{-})\right]\varphi_{\varepsilon}(y,w){\,\mathrm{d}}z_{-}\right)_{+}{\,\mathrm{d}}z_{+}$
@@
with $0\leq\varphi_{\varepsilon}\leq 1$ smooth only depending on $(y,w)$, $\varepsilon>0$, $z_{+}=(x,v,t)\in Q^{+}$ and $z_{-}=(y,w,s)\in Q^{-}$. The second step consists in constructing piecewise affine trajectories following the two vector fields $\mathcal{T}:=\partial_{t}+v\cdot\nabla_{x}$ and $\nabla_{v}$ to connect any $z_{+}\in Q^{+}$ with any $z_{-}\in Q^{-}$:
@@
Note here that a positive time gap between the two cylinders $Q^{-}$ and $Q^{+}$ is used to make sure the intermediate velocity (xy)/(ts)(x-y)/(t-s) remains bounded. We then write $f(z_{+})-f(z_{-})$ as an integral along the chosen trajectory and use ${\mathcal{T}}f+{\mathcal{A}}^{*}{\mathcal{A}}f\leq 0$:
$\int_{Q^{-}}\left[f(z_{+})-f(z_{-})\right]\varphi_{\varepsilon}(y,w){\,\mathrm{d}}z_{-}$ $=\int_{Q^{-}}\int_{\text{trajectory}}{\mathcal{T}}f(\cdots)\varphi_{\varepsilon}(y,w){\,\mathrm{d}}s{\,\mathrm{d}}z_{-}$
$\quad+\int_{Q^{-}}\int_{\text{trajectory}}\nabla_{v}f(\cdots)\varphi_{\varepsilon}(y,w){\,\mathrm{d}}s{\,\mathrm{d}}z_{-}$
$\lesssim\int_{Q^{-}}\int_{\text{trajectory}}\left[\nabla_{v}\cdot\left(A\nabla_{v}f\right)\right](\cdots)\varphi_{\varepsilon}(y,w){\,\mathrm{d}}s{\,\mathrm{d}}z_{-}$
$\quad+\int_{Q^{-}}\int_{\text{trajectory}}\nabla_{v}f(\cdots)\varphi_{\varepsilon}(y,w){\,\mathrm{d}}s{\,\mathrm{d}}z_{-}.$
@@
$(x,v,t)$ $\underset{\nabla_{x}}{\longrightarrow}(x+\varepsilon w,v,t)\underset{\nabla_{v}}{\longrightarrow}\left(x+\varepsilon w,\frac{x+\varepsilon w-y}{t-s},t\right)$
$\underset{{\mathcal{T}}}{\longrightarrow}\left(y,\frac{x+\varepsilon w-y}{t-s},s\right)\underset{\nabla_{v}}{\longrightarrow}(y,w,s).$
This sub-trajectory “noises” the future position variable with the past velocity. This produces a second error term of the form $\varepsilon^{\sigma}\|f\|_{L^{1}_{t,v}W^{\sigma,1}_{x}}$, which can, however, be controlled for a small $\varepsilon$ by the integral regularity $L^{1}_{t,v}W^{\sigma,1}_{x}$ (which can be established for sub-solutions to the Kolmogorov equation, for a small $\sigma>0$). Finally it yields
@@
for a constant $C_{\varepsilon}>0$ depending on $\varepsilon>0$. The two additional errors weaken the Poincaré inequality, but the latter is shown to be sufficient for implementing the next steps of the De Giorgi theory in [8].
@@
(1.23) $\fint_{Q^{+}}\left[f(z_{+})-\langle f\rangle_{Q^{-}}\right]_{+}{\,\mathrm{d}}z_{+}$ $\leq\fint_{Q^{+}}\left(\fint_{Q^{0}}\left[f(z_{+})-f(z_{0})\right]\varphi_{\varepsilon}(x_{0},v_{0}){\,\mathrm{d}}z_{0}\right)_{+}{\,\mathrm{d}}z_{+}$
$\quad+\fint_{Q^{-}}\left(\fint_{Q^{0}}\left[f(z_{0})-f(z_{-})\right]\varphi_{\varepsilon}(x_{0},v_{0}){\,\mathrm{d}}z_{0}\right)_{+}{\,\mathrm{d}}z_{-}$
@@ 1.7. Outline
In Section 2 we provide the explicit construction of the trajectories connecting points in the future to points in the past. These trajectories are then used in Section 3 to prove the Poincaré inequality; they allow to estimate the $L^{1}$ norm in the future of the difference between $f$ and the past average of $f$, or in other words, the left hand side of (1.19). In Section 4, we prove the Harnack inequalities on the basis of this Poincaré inequality.
Acknowledgements
We are grateful to Lukas Niebel and Rico Zacher for stimulating discussions and inspiring ideas on the subject of this article.
@@ 2.1. The base case
[…]
@@
[…]
Figure 3. Construction of the trajectories. The curve $\Gamma_{+}$ connects any point $\Gamma_{+}(0)=z_{+}\in Q^{+}$ to some intermediate point $\Gamma_{+}(1)=z_{0}\in Q^{0}$ along the vector fields $\mathcal{T}$} and $\nabla_{(0)}$.
@@
[…]
Let us denote $\text{{\bf 1}}_{d}$ the $d$-vector of $1$’s, $\text{{\bf 0}}_{d}$ the $d$-vector of 0’s, $\text{Id}_{d}$ the $d\times d$ identity matrix, $\mathbb{O}_{d}$ the $d\times d$ zero block, and
@@
${\mathsf{W}}_{\pm}^{\delta}(s):=\begin{pmatrix}\text{Id}_{d}&\mathbb{O}_{d}\\ \mathbb{O}_{d}&\delta_{\pm}\text{Id}_{d}\end{pmatrix}{\mathsf{W}}(s),\quad\mathsf{Y}_{\pm}=\begin{pmatrix}\mathsf{y}^{(0)}_{\pm}\\ \mathsf{y}^{(1)}_{\pm}\end{pmatrix}:=\begin{pmatrix}x^{(0)}_{0}-x^{(0)}_{\pm}\\ x^{(1)}_{0}-x^{(1)}_{\pm}-\delta_{\pm}x^{(0)}_{\pm}\end{pmatrix}$.
@@
$\begin{pmatrix}X^{(0)}_{\pm}(s)\\ X^{(1)}_{\pm}(s)\end{pmatrix}={\mathsf{W}}_{\pm}^{\delta}(s)\mathsf{M}_{\pm}+\begin{pmatrix}\text{Id}_{d}&\mathbb{O}_{d}\\ \delta_{\pm}s\text{Id}_{d}&\text{Id}_{d}\end{pmatrix}\begin{pmatrix}x_{\pm}^{(0)}\\ x_{\pm}^{(1)}\end{pmatrix}$.
@@
$\begin{pmatrix}X^{(0)}_{\pm}(s)\\ X^{(1)}_{\pm}(s)\end{pmatrix}$ $={\mathsf{W}}_{\pm}^{\delta}(s){\mathsf{W}}_{\pm}^{\delta}(1)^{-1}\mathsf{Y}_{\pm}+\begin{pmatrix}\text{Id}_{d}&\mathbb{O}_{d}\\ \delta_{\pm}s\text{Id}_{d}&\text{Id}_{d}\end{pmatrix}\begin{pmatrix}x_{\pm}^{(0)}\\ x_{\pm}^{(1)}\end{pmatrix}$
@@
$\quad-W^{\delta}_{\pm}(s)\big{[}W^{\delta}_{\pm}(1)\big{]}^{-1}\begin{pmatrix}x^{(0)}_{\pm}\\ x^{(1)}_{\pm}+\delta_{\pm}x^{(0)}_{\pm}\end{pmatrix}+\begin{pmatrix}\text{Id}_{d}&\mathbb{O}_{d}\\ \delta_{\pm}s\text{Id}_{d}&\text{Id}_{d}\end{pmatrix}\begin{pmatrix}x_{\pm}^{(0)}\\ x_{\pm}^{(1)}\end{pmatrix}$
@@
${\mathsf{W}}_{\pm}^{\delta}(1)=\begin{pmatrix}\frac{\text{Id}_{d}}{(1+\alpha_{0})}&\frac{\text{Id}_{d}}{(1+\alpha_{1})}\\ \frac{\delta_{\pm}\text{Id}_{d}}{(1+\alpha_{0})(2+\alpha_{0})}&\frac{\delta_{\pm}\text{Id}_{d}}{(1+\alpha_{1})(2+\alpha_{1})}\end{pmatrix}$,
${\mathsf{W}}_{\pm}^{\delta}(s)^{-1}=\frac{1}{\delta_{\pm}\det{\mathsf{W}}(s)}\begin{pmatrix}\frac{\delta_{\pm}s^{2+\alpha_{1}}\text{Id}_{d}}{(1+\alpha_{1})(2+\alpha_{1})}&\frac{-s^{1+\alpha_{1}}\text{Id}_{d}}{(1+\alpha_{1})}\\ \frac{-\delta_{\pm}s^{(2+\alpha_{0})}\text{Id}_{d}}{(1+\alpha_{0})(2+\alpha_{0})}&\frac{s^{1+\alpha_{0}}\text{Id}_{d}}{(1+\alpha_{0})}\end{pmatrix}$,
@@ 2.2. The general case
[…]
@@
[…]
where $\tilde{\mathsf{B}}_{i,j}:={\mathsf{B}}_{i}{\mathsf{B}}_{i-1}\cdots{\mathsf{B}}_{j}$ for $1\leq j\leq i$, which is a $d_{i}\times d_{j-1}$ block matrix. We denote by $\mathbb{O}$ a zero block matrix with arbitrary size and
@@
(2.12) $\nabla_{(0)}\left(\Phi_{\pm}^{s}\right)^{-1}={\mathsf{W}}_{\pm}^{\delta}(1)\big{[}{\mathsf{W}}_{\pm}^{\delta}(s)\big{]}^{-1}\begin{pmatrix}\text{{\bf 1}}_{d_{0}}\\ \text{{\bf 0}}_{d_{1}}\\ \vdots\\ \text{{\bf 0}}_{d_{\kappa}}\end{pmatrix}$.
@@
Consider the $(\kappa+1)\times(\kappa+1)$ matrix $P(s):=(s^{1+\alpha_{j}+i}/p_{i,j})_{i,j=0}^{\kappa}$ with $p_{i,j}:=(1+\alpha_{j})\cdots(1+i+\alpha_{j})$ and the matrix [^1] The notation $P\otimes\text{Id}_{d_{0}}$ is the classical tensor product between the matrices $P$ and $\text{Id}_{d_{0}}$. $W:=P\otimes\text{Id}_{d_{0}}$. Then their determinants are given by